OpenStudy (anonymous):

5x2/2y3 x 3x/2y2 = ?

razor99 (razor99):

ii mean easy

razor99 (razor99):

thats the answer

OpenStudy (anonymous):

whale.cs.ntou.edu.tw/~chern/courses/doc/tan/PP_Chapter_08.pdf

OpenStudy (anonymous):

\[\frac{5x^2}{2y^3}*\frac{3x}{2y^2}=\frac{15x^3}{4y^5}\] When you multiply variables that have the same base, you add the exponents

OpenStudy (anonymous):

Ya That's What I Thought .

OpenStudy (anonymous):

=5/y3x3x/4y=15x/12y^2=5x/4y^2